Ask
your own question, for FREE!
Mathematics
10 Online
one card is selected from a deck of playing cards. Determine the probability of selecting a card greater than 9 or a black card.
Still Need Help?
Join the QuestionCove community and study together with friends!
The deck has 52 cards. Cards greater than 9 is all 10's, Jacks, Queens, Kings and Aces. That sums up to 20 cards. 4 of these are black. Half of the deck is black. That's 26 cards. So the probability is 20+16=36 out of 52 cards.
Wrong numbers. The probability would be 16+26/52=>42/52
thanks
Of the red cards in the deck, 10 have value greater than 9 (5 diamonds and 5 hearts). There are 26 black cards in the deck. P(selecting a card greater than 9 or a black card) = 10/52 + 26/52 = 36/ 52 = 9/13.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal